Parth (parthkohli):

Challenge: Show that any positive natural number \(a\) can be written as the sum of numbers in the form \(\dfrac{1}{2^n}\) where \(n \in \mathbb{N}\) and a number can be used any number of times, but ALL numbers -- \(\left\{\dfrac{1}{2^n}:n\in\mathbb{N}\right\}\) -- are to be used.

OpenStudy (anonymous):

a= 1/2 +1/2 +1/2 + ...

OpenStudy (anonymous):

I have a solution, but it may just be because I'm interpreting the problem incorrectly.

OpenStudy (anonymous):

where there are 2a numbers of 1/2

OpenStudy (anonymous):

Right. Same thought here, Suaravshakya.

OpenStudy (anonymous):

n=1

Parth (parthkohli):

lol, I'd add something to my problem.

OpenStudy (anonymous):

But there are many solutions. Instead of summing up halves, you could sum up quarters or eighths or sixteenths, or some random combination of those.

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

Well show that any integer cannot be expressed as sum of 1/2^n where n cannotbe repeated

OpenStudy (anonymous):

Same logic will help

OpenStudy (anonymous):

Ah, much more interesting question =) It's fairly intuitive, but to prove it rigorously may be difficult.

Parth (parthkohli):

It's not really rigorous.

OpenStudy (anonymous):

Well, actually Suar, it's not just that any integer cannot be expressed if you disallow n to be repeated. In fact, the sum cannot reach or exceed 1.

Parth (parthkohli):

Actually, if you read the question, I've written "positive number". =/

OpenStudy (anonymous):

Oh hey. Did you edit it? I could have sworn this said integers before.

Parth (parthkohli):

Heh, edited this a minute before you commented that. Any solutions?

OpenStudy (anonymous):

I'm confused by "A number can be used any number of times but all numbers are to be used" How can all numbers be used? If you want all numbers to be used, then I suppose this is not a finite sum, but an infinite sum.

OpenStudy (anonymous):

And I believe this calls not for a proof but for a disproof.

Parth (parthkohli):

So now is the final edit. Take a look at the question.

Parth (parthkohli):

@Hero: Because my homework problems look like\[2x + 3x = 10, \text{ solve for }x.\]

OpenStudy (anonymous):

lololol truefacts

OpenStudy (anonymous):

So assuming you are allowing infinite sums, this is not difficult. \[\Large \sum_{n=1}^{\infty} \frac{1}{2^n} = 1\] Therefore, \[\Large a\sum_{n=1}^{\infty} \frac{1}{2^n} = a\]

OpenStudy (anonymous):

\[a=\frac{1}{2}k+0\left(1+\sum_{i >1}\frac{1}{2^i}\right), \quad \text{ where } k: a=\frac{1}{2}k.\]

OpenStudy (anonymous):

@SmoothMath Infinite sums are stated explicitly to be used in the definition of all numbers. Your answer is on par with mine in terms of Smart Alec material. This is a compliment.

Parth (parthkohli):

Yes! Just a little note: Use\[\sum_{n =1}^{\infty} \dfrac{1}{2^n} + \sum_{n =1}^{\infty} \dfrac{1}{2^n}+ \sum_{n =1}^{\infty} \dfrac{1}{2^n} \cdots(a \ \text{times}) = a\]because you aren't allowed to use the number \(a\) since it is strictly not in the form shown above...

OpenStudy (anonymous):

\[\sum_{n =1}^{\infty} \dfrac{1}{2^n} + \sum_{n =1}^{\infty} \dfrac{1}{2^n}+ \sum_{n =1}^{\infty} \dfrac{1}{2^n} \cdots(a \ \text{times}) = a\sum_{n \ge 1}2^{-n}\]

OpenStudy (anonymous):

It is a good thing to point out, and I almost did fix it but was too lazy. Another way to satisfy the requirement is by the property of sums: \(\Large a\sum_{n=1}^{\infty} \frac{1}{2^n} = \sum_{n=1}^{\infty} a*\frac{1}{2^n}\)

OpenStudy (anonymous):

Ah, and Limitless said the same thing =)

OpenStudy (anonymous):

@ParthKohli Why don't you prove that this works for _only_ \(x=\dfrac{1}{2}\) in \[\sum_{n \ge 1}x^{n}?\]

OpenStudy (anonymous):

or something similar.

OpenStudy (anonymous):

@SmoothMath To be rough with you, Smooth, I simply demonstrated that I am a bit of a smart alec for the second time in this thread in an attempt to be humorous.

Parth (parthkohli):

@Limitless: What does?

OpenStudy (anonymous):

@ParthKohli Show that \[a=a\sum_{n \ge 1}x^n \iff x=\frac{1}{2}.\]

Parth (parthkohli):

Hmm, that's just another example of infinite geometric series.

Parth (parthkohli):

\[S_n = \dfrac{a}{1 - r} ~ \iff r > |1| \]

Parth (parthkohli):

Oops, I meant\[r < |1|\]

OpenStudy (anonymous):

That's close. Be specific.

Parth (parthkohli):

OK, do you need the derivation of it?